When Jorge took his place in the marching band, he noticed that he was 10th from the front, 7th from the back, 3rd from the left, and 8th from the right in a rectangular formation. How many people are in the marching band?
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

When Jorge took his place in the marching band, he noticed that he was 10th from the front, 7th from the back, 3rd from the left, and 8th from the right in a rectangular formation.
How many people are in the marching band?
:
Vertically: 10 + 7 = 17
Horizontally: 3 + 8 = 11
Therefore
17 * 11 = 187 people in the band

Answer by ikleyn(52781)

.
The column length is 9 + 1 + 6 = 16 people.

The row    length is 2 + 1 + 7 = 10 people.


In all, there are 16*10 = 160 people in the band.